# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

### v1

@@ A.2 Properties of Sobolev Spaces
Proposition A.4 (Bound of Potential Energy [1, Proposition 6]).
$\exists C>0$, if $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, then $\left|\int\rho\cdot V_{\rho}\right|\leq C\int|\rho|^{\frac{4}{3}}\cdot\left(\int|\rho|\right)^{\frac{2}{3}}$, where $V_{\rho}(x)=$ $\int\frac{\rho(y)}{|x-y|}d^{3}y$.
@@
Proposition A.5 (Bound of Potential [1, Proposition 5] [7, Proposition A.11]).
Suppose $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$. If $1<p\leq\frac{3}{2}$, then $\forall r\in\left(3,\frac{3p}{3-2p}\right),V_{\rho}\in L^{r}\left(\mathbb{R}^{3}\right)$, and $\exists 0<b_{r}<1,0<c_{r}<1,C>0$, such that
@@
It turns out we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int\frac{y_{j}\rho(x-y)}{|y|^{3}}d^{3}y$ not only in the distribution sense but also in the classical sense:
Proposition A.6 (Differentiability of Potential [1, Proposition 7] [7, Proposition A.13]).
If $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$ for some $p>3$, then $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)$ is continuously differentiable and the weak derivative coincides with the classical one for all $x\in\mathbb{R}^{3}$.
To prove the low semicontinuity of the energy, we hope to connect the weak convergence of densities and strong convergence of potentials.
Proposition A.7 (Compactness of Convolution (Potential) Operator with $3<q<\infty$ [7, Corollary A.15]).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{N}\right)$. Let F=GBF=G*B, where $B$ is a bounded set in $L^{q}\left(\mathbb{R}^{N}\right)$ with $3<q<\infty$. Then let $\Omega\subset\mathbb{R}^{3}$ is open, bounded and of class $C^{1},\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$. In particular, let $V_{\rho}(x)=\int\frac{\rho(y)}{|x-y|}d^{3}y,\forall 1\leq r\leq\infty$, the map $\mathcal{V}:\rho\rightarrow V_{\rho}$ from $L^{q}(\Omega)$ to $L^{r}(\Omega)$ is compact.
@@
By Proposition A.6 we know $F\subset W^{1,\infty}\left(\mathbb{R}^{3}\right)$. Follow the proof in [7] in detail one can see the norm $\|\cdot\|_{W^{1,\infty}}$ is bounded uniformly. The result comes from Rellich-Kondrachov Theorem [3, Theorem 9.16]. ∎
@@
Proposition A.8 (Compactness of Convolution (Potential) Operator with $q\geq 1$ [7, Corollary A.16]).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{3}\right)$. Let F=GBF=G*B, where $B$ is a bounded set in $L^{q}\left(\mathbb{R}^{3}\right)$ with $q\geq 1$. Then let $\Omega\subset\mathbb{R}^{3}$ is open, bounded and of class $C^{1}$.
@@
    If $1\leq q<\frac{3}{2}$, then $\forall 1\leq r<p$, where $\frac{1}{p}=\frac{1}{q}-\frac{2}{3},\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
@@
    If $q=\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\forall 1\leq r<\infty,\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
@@
    If $q>\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$.
@@
The proof is similar to the proof of Proposition A.5 in [7]. For Parts (2)(3), the proof is essentially the same. For part (1), instead of using Interpolation Inequality, one can use Hardy-Littlewood-Sobolev Inequality Proposition A.3 to show $F$ is bonded in $W^{1,s}$ for some $s$, and then use Rellich-Kondrachov Theorem [3, Theorem 9.16] as well . ∎
Remark A.9.
There is another compactness result about the convolution operator in [3, Corollary 4.28], but it’s not valid for potential since $G(x)=\frac{1}{|x|}$ is not in $L^{p}$.
@@ Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez (master’s thesis advisor) and Dimitri Cobb (second advisor). Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.

### v2

@@ Appendix A Properties of Sobolev Spaces
Proposition A.2 (Bound of Potential Energy [4, Proposition 6]).
$\exists C>0$, if $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, let $V_{\rho}(x)=$ $\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$, then
$\left|\int_{\mathbb{R}^{3}}\rho\cdot V_{\rho}\,dx\right|\leq C\int_{\mathbb{R}^{3}}|\rho|^{\frac{4}{3}}\,dx\cdot\left(\int_{\mathbb{R}^{3}}|\rho|\,dx\right)^{\frac{2}{3}}$
@@
Proposition A.3 (Bound of Potential [4, Proposition 5] [17, Proposition B.1]).
Suppose $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$. If $1<p\leq\frac{3}{2}$, then $\forall r\in\left(3,\frac{3p}{3-2p}\right)$, we have $V_{\rho}\in L^{r}\left(\mathbb{R}^{3}\right)$, and $\exists 0<b_{r}<1,0<c_{r}<1,C>0$, such that
@@
As discussed in [17, Appendix B], it turns out $h(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}\,dy$ is the weak derivative of $V_{\rho}(x)$. Moreover, for $p>3$, we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}\,dy$ not only in the distribution sense but also in the classical sense:
Proposition A.4 (Differentiability of Potential [4, Section 3] [17, Proposition B.3]).
If $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$ for some $p>3$, then $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)$ is continuously differentiable, and the weak derivative coincides with the classical one for all $x\in\mathbb{R}^{3}$.
To prove the low semicontinuity of the energy, we hope to connect weak convergence of densities and strong convergence of potentials.
Proposition A.5 (Compactness of Convolution (Potential) Operator with $3<q<\infty$).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{N}\right)$. Let F=GBF=G*B, where $B$ is a bounded set in $L^{q}\left(\mathbb{R}^{N}\right)$ with $3<q<\infty$. Let $\Omega\subset\mathbb{R}^{3}$ be open, bounded and of class $C^{1}$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$. In particular, let $V_{\rho}(x)=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$, then for all $r\in[1,\infty]$, the map $\mathcal{V}:\rho\rightarrow V_{\rho}$ from $L^{q}(\Omega)$ to $L^{r}(\Omega)$ is compact.
@@
By Proposition A.4 we know $F\subset W^{1,\infty}\left(\mathbb{R}^{3}\right)$. Following its proof in [17] in detail one can see the norm $\|\cdot\|_{W^{1,\infty}}$ is bounded uniformly. The result comes from Rellich-Kondrachov Theorem [8, Theorem 9.16]. ∎
@@
Proposition A.6 (Compactness of Convolution (Potential) Operator with $q\geq 1$).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{3}\right)$. Let F=GBF=G*B, where $B$ is bounded both in $L^{1}\left(\mathbb{R}^{3}\right)$ and $L^{q}\left(\mathbb{R}^{3}\right)$ with $q>1$. Let $\Omega\subset\mathbb{R}^{3}$ be open, bounded and of class $C^{1}$, then
@@
    if $1<q<\frac{3}{2}$, then $\forall 1\leq r<p$, where $\frac{1}{p}=\frac{1}{q}-\frac{2}{3},\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$;
@@
    if $q=\frac{3}{2}$, then $\forall 1\leq r<\infty,\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$;
@@
    if $q>\frac{3}{2}$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$.
@@
We first establish the boundedness of $F$ in $W^{1,s}$ for some appropriate $s$. Similar to the proof of [17, Proposition B.1] (cf. Proposition A.3 above), and noting that $K(y)=\frac{y_{j}}{|y|^{3}}$ is in $L^{p}_{\text{loc}}(\mathbb{R}^{3})$ for $p<3/2$ and belongs to $L^{p}(\mathbb{R}^{3}\setminus B_{R}(0))$ for $p>3/2$, we can show there exists an appropriate $s>0$, such that $\forall\rho\in B$, $V_{\rho}(x)$ and $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}\,dy$ are bounded in $L^{s}$. Since $B$ is bounded in $L^{1}$ and $L^{q}$, the bound of $V_{\rho}$ in $W^{1,s}$ is uniform for all $\rho\in B$. Once the boundedness of $F$ in $W^{1,s}$ is established, we apply the Rellich–Kondrachov theorem [8, Theorem 9.16] to conclude the proof.
@@ Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez and Dimitri Cobb for their continued advice and support since the author’s time in Bonn, as well as Christof Sparber and Mimi Dai for their comments and support during the author’s Ph.D. studies at the University of Illinois Chicago. Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.